40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Aug 23 00:55:13 2011, Jenne, Update, General, IFO ready for doors 
    Reply  Wed Aug 31 14:50:18 2011, kiwamu, Update, Photos, pictures of OSEMs 
Message ID: 5280     Entry time: Tue Aug 23 00:55:13 2011     Reply to this: 5329
Author: Jenne 
Type: Update 
Category: General 
Subject: IFO ready for doors 

[Kiwamu, Jenne]

After the IFO was aligned in air one final time, we tapped on a few OSEMs until we were happy with all of the centering of all of the optics' OSEMs.  All are within 0.05 of their halfway values, with the exception of one each on MC1 and MC3, one of which is within 0.06, and the other 0.08.  Because of the realignment pain of dealing with MC OSEMs, we elected to leave these alone.  Also, since we obviously didn't open the MC2 tank, we don't know how they are, although the numbers look reasonable. 

Also, we took photos (to be posted on Picasa in a day or two) of all the main IFO magnet-in-OSEM centering, as best we could.  SRM, BS, PRM all caused trouble, due to their tight optical layouts.  We got what we could.  Various people have been looking at these for the past 2 weeks, and I think they're all fine, even if we didn't get stellar photos.

We are now prepared for pumping.  For real this time.

 

ELOG V3.1.3-